Trace the curve x=a(theta + sin theta), y = a(1-cos theta). State the properties you use for tracing it also.


1
Expert's answer
2022-01-28T13:19:29-0500


We use the following properties for tracing the curve. We have the Cartesian curve defined by the parametric equations "x=f(\theta), y = g(\theta)."

Since "\ud835\udc66" is a periodic function of "\theta" with period "2\pi," it is sufficient to trace the curve for "\theta\in[0, 2\pi]." For "\theta\in[0, 2\pi], x" and "y" are well defined.

Trace the curve "x=a(\theta + \sin \theta), y = a(1-\cos \theta), 0\leq\theta\leq 2\pi, a>0."

1. Symmetry


"f(-\theta)=a(-\theta+\sin(-\theta))=-f(\theta),"

"g(-\theta)=a(1-\cos(-\theta))=g(\theta)"

Therefore, the curve is symmetrical about the 𝑦 −axis. Curve is not symmetrical about 𝑦- axis.


2. Origin

"(0, 0), x=f(\theta)=a(\theta + \sin \theta)=0,"


"y=g(\theta) = a(1-\cos \theta)=0"

"\begin{cases}
 \sin\theta=-\theta \\
 \cos\theta=1
\end{cases}=> \begin{cases}
 \sin\theta=-\theta \\
 \theta=0\ \ or\ \ \theta=2\pi
\end{cases}=>\theta=0"

A curve passes through the origin.

Derivatives:


"\dfrac{dy}{dx}=\dfrac{\dfrac{dy}{d\theta}}{\dfrac{dx}{d\theta}}=\dfrac{a\sin \theta}{a(1+\cos \theta)}=\dfrac{\sin \theta}{1+\cos \theta}=\tan(\theta/2)"

At "\theta=0, \dfrac{dy}{dx}=0." Tangent to the curve at "\theta=0" is "x" -axis.

At "\theta=\pi, \dfrac{dy}{dx}=\infin." Tangent to the curve at "\theta=\pi" is perpendicular to "x" -axis.


3. Intercepts

Intersection with "x"-axis: The points of intersection of the curve with the "x"-axis are given by the roots of "g(\theta)=a(1-\cos \theta)=0, 0\leq \theta \leq 2\pi, a>0."


"\cos \theta=1"

"\theta=0\ or\ \theta=2\pi"

"x=0\ or\ x=2\pi a"


Then "(x, y)=(0, 0), (x, y)=(2\pi a, 0)."


Intersection with "y"-axis: The points of intersection of the curve with the "y"-axis are given by the roots of "f(\theta)=a(\theta+\sin \theta)=0, 0\leq \theta \leq 2\pi, a>0."

"\theta+\sin \theta=0"

"\theta=0"

"y=0"


Then "(x, y)=(0, 0)."


4. Asymptotes

"x=a(\theta+\sin \theta), y=a(1-\cos \theta)."

There is no vertical asymptote.

There is no horizontal asymptote.

There is no oblique asymptote.


5. Regions where no Part of the curve lies

"x=a(\theta+\sin \theta), y=a(1-\cos \theta)."

Note that "y\geq 0." Entire curve lies above the "y" -axis "(0\leq y\leq 2a)."


6. First derivative

"x=a(\theta+\sin \theta), y=a(1-\cos \theta)."


"\dfrac{dy}{dx}=\dfrac{\dfrac{dy}{d\theta}}{\dfrac{dx}{d\theta}}=\dfrac{a\sin \theta}{a(1+\cos \theta)}=\dfrac{\sin \theta}{1+\cos \theta}=\tan(\theta/2)"

At "\theta=0, \dfrac{dy}{dx}=0." Tangent to the curve at "\theta=0" is "x" -axis.

At "\theta=\pi, \dfrac{dy}{dx}=\infin." Tangent to the curve at "\theta=\pi" is perpendicular to "x" -axis.

For "0<\theta<\pi,\dfrac{dy}{dx}>0."

The function "y(x)" is increasing in this interval.

For "\pi<\theta<2\pi,\dfrac{dy}{dx}<0."

The function "y(x)" is decreasing in this interval.


8. Second derivative

"x=a(\theta+\sin \theta), y=a(1-\cos \theta)."


"\dfrac{d^2y}{dx^2}=\dfrac{\dfrac{d}{d\theta}(\dfrac{dy}{dx})}{\dfrac{dx}{d\theta}}=\dfrac{\dfrac{a}{2\cos^2(\theta/2)}}{a(1+\cos \theta)}=\dfrac{1}{4\cos^4(\theta/2)}"


For "0<\theta<2\pi, \dfrac{d^2y}{dx^2}>0," => y is concave upward.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 \theta & 0 & \pi/2 & \pi & 3\pi/2 & 2\pi \\ \hline
 x & 0 & a(\pi/2+1) & a\pi & a(3\pi/2-1) & 2a\pi\\
 \hdashline
 y & 0 & a & 2a & a & 0\\
 \hdashline
 dy/dx & 0 & 1 & \infin & -1 & 0 \\
 \hdashline
\end{array}"